# Enhancing Video Large Language Models with Structured Multi-Video Collaborative Reasoning (early version)

### v1

@@ 5 Conclusion
This paper introduces a pioneering framework that enhances video large language models through structured multi-video collaborative reasoning. We first design the Video Structuring Module that models video as a spatio-temporal graph. Then, The Graph Fusion Module integrates related video information into enhanced graph tokens, which are then combined with visual and textual tokens in a multi-video structured prompt for input to the language model. Extensive experiments demonstrate the effectiveness and robustness of our approach to comprehending complex video content and answering queries accurately. We hope our work can provide insights for reliable video understanding and stimulate more research interest.

### v2

@@ 5 Conclusion
In this work, we introduce a pioneering framework that enhances video large language models through structured multi-video collaborative reasoning. We first design the Video Structuring Module that models video as a spatio-temporal graph. Then, the Graph Fusion Module integrates related video information into enhanced graph tokens, which are then combined with visual and textual tokens in a multi-video structured prompt for input to the language model. Extensive experiments demonstrate the effectiveness and robustness of our approach in comprehending complex video content and answering queries accurately. We hope our work can provide insights for reliable video understanding and stimulate more research interest.
6 Statements and Declarations
6.1 Data Availability Statement
The original video source and corresponding annotations are legally accessible through the publicly available open-source dataset [zhang2024videoinstructiontuningsynthetic]. Furthermore, we plan to publicly release both our code for the data processing pipeline and the training dataset utilized for GFM training in the near future.
6.2 Acknowledgment and Competing Interests
The paper is supported in part by the National Natural Science Foundation of China (No. 62325109, 62561160155, U21B2013), and in part by the Shanghai ‘The Belt and Road’ Young Scholar Exchange Grant (24510742000).